Suppose that the heat flow out of your 20°C home in the winter is 7

kW. If the temperature outside is -15° C, how much power P would an

ideal heat pump require to keep your inside temperature at 20° C?


1
Expert's answer
2022-04-14T15:45:52-0400

"\frac{T_2}{T_1}=\frac{P_2}{P_1},"

"P_1=P_2\frac{T_1}{T_2}=8~kW."